OpenStudy (anonymous):

Mrs. Hernanadez is covering the top of a kitchen shelf that is 26 inches long and 15 inches wide. She incorrectly measures the length to be 25 inches long. How will this error affect her calculations?

OpenStudy (anonymous):

area = 26*15 = 390 calculated = 25*15 = 375 , error is their difference

OpenStudy (calculusxy):

It is going to affect the area of the kitchen shelf since you are having different factors of multiplying to get the area. We know that 25*15=375 And 26*15=390 So it w l have difference of 15 square inches

OpenStudy (anonymous):

thank you so much guys!!!!

OpenStudy (calculusxy):

Your welcome :)

OpenStudy (anonymous):

ur welcome :D